OpenStudy (anonymous):

Factor completely 9xy + 36x − 10y – 40.

OpenStudy (anonymous):

@slinley

OpenStudy (anonymous):

@iYuko

iYuko (iyuko):

9xy + 36x - 10y - 40 9x(y + 4) - 10(y + 4) = (9x - 10)(y + 4) See? Do you understand now?

iYuko (iyuko):

Leah?

OpenStudy (anonymous):

no

iYuko (iyuko):

( 9x -10 ) ( 4 + y) equation is factred x= 10/9 and y= - 4

iYuko (iyuko):

Factored*

OpenStudy (anonymous):

that is not my answer choices

iYuko (iyuko):

What r the choices?

OpenStudy (anonymous):

im going to type them gimme a min

OpenStudy (anonymous):

a.)prime

OpenStudy (anonymous):

(9xy+36x) - (10y+40) is the equation simplified. Explanation: Everything staid the same except for the -40, it is now + 40 because of the other minus sign

iYuko (iyuko):

(9x-10)(y+4)

OpenStudy (anonymous):

b.) (y-4)(9x-10)

iYuko (iyuko):

Its b

OpenStudy (anonymous):

@iYuko No direct answers please

OpenStudy (anonymous):

ok can u help me with more

iYuko (iyuko):

Kitten lover I helped her then gave her the answer so idk what your meaning of "direct" is

iYuko (iyuko):

Leah close this question then tag me to the next one